OpenStudy (anonymous):

What is the center of the circle with the equation (x+1^)2 + (y-3)^2= 9?

OpenStudy (anonymous):

well the center of your circle would be (-1,3) because the equation for a circle is (x-a)^2+(y-b)^2=r^2 so because your equation is (x+1)^1+(y-3)^2=9 your center point would be (-1,3) and your radius is 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome